OpenStudy (anonymous):

3√ = x+2 solve radical equation

OpenStudy (bahrom7893):

that doesn't make sense.

OpenStudy (saifoo.khan):

\[\sqrt3 = x+2\]?

OpenStudy (anonymous):

oh wait 3√x = x+2

OpenStudy (bahrom7893):

9x = x^2 + 4x + 4 x^2 -5x + 4 = 0

OpenStudy (bahrom7893):

x = 1, or 4

OpenStudy (anonymous):

how did u get the 9x part

OpenStudy (bahrom7893):

square both sides